The length of a rectangle is 2 more than twice its width. If the area of the rectangle is 24 sq. meters, find its dimension
Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

2 more than twice its width. If the area of the rectangle is 24
Area of the rectangle:
L*W = A
:
L = 2W
Therefore:
2W*W = 24
2W^2 = 24
W^2 = 12
W = sqrt12
or
W = -sqrt12
W = 3.464
or
W = -3.464
Since we are looking for a positive number, the answer is:
W = 3.464
--------------------
Check:
L*W = 24
2W*W = 24
2(3.464)*3.464 = 24
6.928*3.464 = 24 Correct.
John

Answer by MathTherapy(10552)


The length of a rectangle is 2 more than twice its width. If the area of the rectangle is 24 sq. meters, find its dimension